Maribel has $1.25 in her pocket. The money is in quarters and dimes. There are a total of 8 coins. How many quarters and dimes does Maribel have in her pocket?
Answer by mananth(16949)

Maribel has $1.25 in her pocket. The money is in quarters and dimes. There are a total of 8 coins. How many quarters and dimes does Maribel have in her pocket?
Let the number of quarter coins be x
The number of dimes will be 8-x
The value of quarter will be 5x
The value of dimes will be 10*(8-x)
Their sum = 125 cents
25x+10*(8-x)=125
25x+80-10x=125
25x-10x=125-80
15x=45
x=3
quarters will be 3
Dimes will be 5